OpenStudy (anonymous):

Determine the equation of a line. a)passes through the points (2,5) and (4,9). ? anyone know

OpenStudy (anonymous):

Calculate slope which is 2

OpenStudy (anonymous):

2x-y=-1

OpenStudy (anonymous):

y=2x+1

OpenStudy (anonymous):

use the point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

that should be the right answer